# Macroeconomic Spillovers of Weather Shocks across U.S. States

### v2

@@ 2.2 State-level economic variables
State-level economic activity. Our aim is to capture the impact of weather-related disasters on broad measures of economic activity that allow assessing the presence of spillover effects. Unfortunately, data on Gross State Product is only available at a quarterly frequency and only from 2005, resulting in a relatively small number of available observations. Moreover, working with monthly data appears preferable in this context, since the duration of extreme weather events is generally shorter than one month, and so using lower-frequency data may lead to temporal aggregation bias. The proxy of economic activity developed by Baumeister et al., (2024), referred to as weekly Economic Conditions Indicators (ECI), is suitable for our purposes. This is a composite indicator based on mixed-frequency dynamic factor models covering multiple dimensions of state-level economic activity (e.g., mobility, labor market, real activity, expectations measures, financial and households indices). Since the ECI is scaled to match four-quarter growth rates of U.S. real GDP, and we are interested in measuring the economic effects of weather shocks at a monthly frequency, we take the monthly average of the weekly index and label it Monthly Economic Conditions Indicator (MECI) hereafter. Unlike other indicators of economic activity available at the state level on a monthly basis - such as employment or hours worked - the MECI includes a broad set of both standard and non-standard variables (e.g., electricity consumption, oil rig counts, vehicle miles traveled) that allow for a more comprehensive measure, enabling a better study of any spillover effects. See Bokun et al., (2023) for an overview of state-specific variables for the U.S. [^6] In the online Appendix, we rely on the number of hours worked in manufacturing as an alternative dependent variable in the GVAR, in place of MECI. As expected, being a less comprehensive measure of economic activity, the impact of weather shocks on hours is stronger in the region hit by the shock, while spillover effects in other states are often negligible.
@@
Note that we are interested in studying the effects of extreme weather, therefore it is important to assess the presence of seasonality in the MECI. The individual components of the ECI are seasonally adjusted when appropriate (refer to Baumeister et al., , 2024, for details on seasonal adjustment). Nevertheless, we further check that there are no seasonal effects left in the monthly version of the index by estimating state-specific autoregressive models augmented with month-of-the-year dummy variables. The $F$-tests indicate for all states, but North Carolina, month-of-the-year dummies are jointly non-significant, and therefore seasonal factors are not expected to affect our results.
@@
Interstate trade flows. In GVAR models, economic interrelationships between geographical entities are most commonly measured using trade flows. To proxy bilateral trade flows between U.S. states, we use data from the latest release of the Commodity Flow Survey (CFS) in 2017. The CFS is jointly elaborated by the Bureau of Transportation Statistics (U.S. Department of Transportation) and the U.S. Census Bureau (U.S. Department of Commerce) every five years as part of the Economic Census. This survey is the primary source of data on domestic freight shipments by establishments in mining, manufacturing, wholesale, and selected retail and services trade industries (i.e., electronic shopping and mail-order houses, fuel dealers, publishers) located in the U.S. We use these data to construct weight matrices that summarize network links between states. In particular, to calculate the weight of state $j$ in the network of state $i$, we use the sum of bilateral imports and exports between the two states (i.e., volumes of CFS shipments of goods from $j$ to $i$ and from $i$ to $j$, respectively), as a ratio of total trade of state $i$ (imports into $i$ plus exports from $i$) with all other states in the U.S.
@@
Notes: The left panel shows the classification of U.S. states into nine climate regions, provided by the National Oceanic and Atmospheric Administration (source: https://www.ncei.noaa.gov/access/monitoring/reference-maps/us-climate-regions). Hawaii and Alaska are not considered in this classification but are kept in the analysis. The right panel displays the network visualization of the bilateral trade flows between U.S. states.
On the left panel of Figure 5 we show the nine climate regions according to the classification provided by the National Oceanic and Atmospheric Administration (Karl and Koss, , 1984). The nine climate regions are: Northeast (NE), Southeast (SE), South (S), Upper Midwest (UMW), Ohio Valley (OV), Northern Rockies and Plains (NP), Southwest (SW), West (W), and Northwest (NW).
A network representation of the interstate trade flows, instead, is provided in the right panel of Figure 5. Visual clutter is reduced using CFS data to determine, for each state, the three most important trade partners and representing the network as an undirected graph. Nodes are proportional to the size of the local economy, as proxied by state employment in 2017 sourced from the FRED-SD database (Bokun et al., , 2023), and have the same color as the climate region to which they belong, as shown in the left panel of the figure. The network reflects the underlying spatial proximity of states as well as their centrality to the U.S. economy as a whole. For instance, we can observe a cluster of northeastern economies tightly interconnected on the right portion of the graph. Moreover, it is evident that New York, Texas, Florida, and California, not only represent the largest economies in terms of employment, but are also heavily connected with several other states, thus playing a pivotal role in the entire U.S. economy.
@@
We construct a GVAR model for the U.S. economy by aggregating state-specific autoregressive models ($ARX^{\star}$) that include rest-of-U.S. (RoUS) exogenous variables, along with weather shocks. RoUS aggregate variables are calculated as weighted averages across all other states, using inter-state trade shares as weights. Since RoUS aggregates are treated as weakly exogenous in each $ARX^{\star}$ model, estimation can be performed at the state level, thus avoiding overparametrization problems. Let $y_{it}$ denote an indicator of economic activity for state $i$ at time $t$ (i.e., MECI), and $s_{it}$ be our state-level weather-shock proxy. For ease of notation, in the following discussion we consider a model with one lag. The $ARX^{\star}$ for state $i$ is then given by:
@@
$\begin{pmatrix}y_{it}\\ y_{it}^{\ast}\end{pmatrix}=\mathbf{W}_{i}\begin{pmatrix}y_{1t}&y_{2t}&\bf\cdots&y_{Nt}\end{pmatrix}^{\prime}=\mathbf{W}_{i}\mathbf{y}_{t}$ (4)
@@
Stacking all $ARX^{\star}$ models, we get a country-wide model:
@@
$\mathbf{G}=\begin{pmatrix}\left(1,-\gamma_{01}\right)\mathbf{W}_{1}\\ \left(1,-\gamma_{02}\right)\mathbf{W}_{2}\\ \vdots\\ \left(1,-\gamma_{0N}\right)\mathbf{W}_{N}\end{pmatrix}\hskip 35.0pt\mathbf{H}=\begin{pmatrix}\left(\beta_{1},\gamma_{11}\right)\mathbf{W}_{1}\\ \left(\beta_{2},\gamma_{12}\right)\mathbf{W}_{2}\\ \vdots\\ \left(\beta_{N},\gamma_{1N}\right)\mathbf{W}_{N}\end{pmatrix}\hskip 35.0pt\boldsymbol{\Theta}=\begin{pmatrix}\theta_{1}&0&\dots&0\\ 0&\theta_{2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\theta_{N}\end{pmatrix}$ (6)
while $\boldsymbol{\alpha}$, $\mathbf{s}_{t}$ and $\mathbf{u}_{t}$ are vectors of stacked constants, weather shocks and errors, respectively. By inverting the matrix $\mathbf{G}$, we obtain a reduced-form VAR for the entire U.S. economy:
@@ 4 Empirical Results
We estimate our monthly GVAR from January 1990 to December 2019. We exclude the subsequent period to prevent our results from being biased by the COVID-19 pandemic. We rely on a GVAR model with 2 lags as suggested by minimization of the Bayesian information criterion, which selects at most 2 lags for almost all states. [^7] For robustness, we have also estimated the model using, alternatively, lag orders of 1 and 3. The results remain very similar in both cases. Furthermore, in consideration of the monthly frequency of the data, we have also estimated the model with 12 lags. Results are qualitatively similar also in this case. We derive our baseline results using the MECI as a measure of local economic activity. Unit root tests indicate that the level of the MECI is non-stationary or near unit root for a number of states, and stationary for others. Therefore, we transform the MECI indices by taking their first differences and estimate each ARX* model (3) by OLS. [^8] Since not all MECI indices are non-stationary, we cannot represent our U.S. GVAR model as a cointegration model as in Pesaran et al., (2004). Moreover, we have tried estimating the GVAR on the levels of the MECI, treating them as stationary, but the model turns out to have explosive roots in this case. Reliance on first differences of MECI implies that our model does not have an error correction form. For the same reason, we cannot rely on the usual likelihood ratio (LR) test for weak exogeneity of $y_{it}^{*}$ used in the cointegrated GVAR literature, which tests the significance of loading coefficients in error-correction country-specific models (see, e.g., Dées et al., 2007a ). However, we perform Granger causality tests on $y_{it}$ and $y_{it}^{*}$ in each state. The tests indicate that, for the vast majority of states, including the ones with the largest economies, the domestic variable $y_{it}$ does not Granger cause the rest-of-U.S. variable $y_{it}^{*}$, while $y_{it}^{*}$ Granger causes $y_{it}$, as expected.
@@
We assess the direct and indirect macroeconomic effects of weather shocks by means of impulse response functions (IRFs) estimated on a sample that runs from April 1990 through December 2019, for a total of 357 observations (i.e., after taking first differences of MECI and considering two lags). While we carry out the impulse-response analysis at the state level, in what follows, to summarize results, we aggregate the estimated individual IRFs into those for nine U.S. climate regions, as previously shown in Figure 5. Note that while Hawaii and Alaska are not considered in this classification, we do retain them in the GVAR analysis.
@@
When simulating a weather shock, we assume that all states in a region are contemporaneously hit. As a matter of fact, Figure 3 highlights that major weather events such as the Storm of the Century in 1993, the Blizzard of 1996, Hurricane Katrina in 2005 and the Super Outbreak in 2011 have triggered emergency declarations in a very large portion of the U.S. This is thus an extremely adverse but realistic scenario, that allows us to interpret the results as “upper bounds” of cross-regional spillover effects.
4.1 Main results
@@
We assess the responses of MECI to a region-wide weather shock in Figures 6 and 7. While the GVAR model is estimated using the first differences of MECI, we cumulate the Impulse Response Functions (IRFs) to obtain the level responses of the index, which is scaled to match the four-quarter percent growth rate of U.S. GDP. Figure 6 shows the IRFs for a subset of regions (i.e., NE, SE, S, and W), while Figure 7 reports the IRFs for all regions one year after the shock, which seems a reasonable time frame to assess the magnitude of macroeconomic effects induced by weather disasters. We compute means and the corresponding 10th and 90th percentiles using a bootstrap procedure for GVAR models (see Dées et al., 2007a, ; Dées et al., 2007b, , for details). [^9] The bootstrap procedure is non-parametric, so we do not need to make any assumption on the distribution of errors $u_{it}$ in equation (3).
@@
Notes: Responses of the Monthly Economic Conditions Indicator (MECI) to a weather shock hitting 100% of counties in the shocked region. The horizontal axis shows months, the vertical axis the responses in percentage points. Blue solid lines indicate bootstrap means and the red area the corresponding 10th and 90th percentile. Four climate regions are considered: Northeast (NE), Southeast (SE), South (S), and West (W). In titles, the first region indicates where the shock originates, while the second one is where it propagates.
@@
Notes: The figure displays IRF bootstrap means along with their corresponding 10th and 90th percentile. The regions affected by shocks are specified on the x-axis, while the title identifies the origin of the shocks.
Remarkably, responses of real economic activity are generally negative, both within the shocked region and in regions not directly affected by the weather shock. Except for Western regions (W, NW, and SW), all regions generate significant spillovers. The IRFs generally converge within approximately one year after the shock. Because of the unit roots/near unit-root behavior of MECI, the model predicts permanent effects, but these may be simply considered as approximations of highly persistent effects, in line with the findings of previous literature on the effects of extreme climatic events (Batten, , 2018; Hsiang and Jina, , 2014).
Figure 7 reports the mean bootstrap responses with the corresponding 10th and 90th percentiles one year after the shock. The largest spillovers (in absolute values) are generated by the South (S), Southeast (SE) and Ohio Valley (OV) regions. Concerning the magnitude of the effects, the mean responses after one year are generally in the range from -1.3 to 0. This indicates, approximately, a drop of up to 1.3% in year-on-year growth rates of economic activity, in response to the region-wide weather shock. The single largest IRF (in absolute values) is represented by the response of the Upper Midwest (UMW) region to a weather shock in the South (S). [^10] To account for potentially non-linear effects of weather shocks, we have estimated the model using squared weather shocks and thresholds for the number of counties hit. The results, not reported here to save space, are qualitatively similar to those reported in Figure 7, and are available from the authors upon request.
4.2 The role of economic linkages and spatial adjacency
The GVAR model allows to estimate spillover effects, taking into account the economic interdependence between states. In this section, we assess the importance of explicitly considering economic linkages using interstate trade flows. To this aim, we re-estimate the GVAR model using alternative weight matrices $\mathbf{W}_{i}$ based on spatial adjacency.

### v3

@@ 2.2 State-level economic variables
State-level economic activity. Our aim is to capture the impact of weather-related disasters on broad measures of economic activity that allow assessing the presence of spillover effects. Unfortunately, data on Gross State Product is only available at a quarterly frequency from 2005, resulting in a relatively small number of available observations. Moreover, working with monthly data appears preferable in this context, since the duration of extreme weather events is generally shorter than one month, and so using lower-frequency data may lead to temporal aggregation bias. The proxy of economic activity developed by Baumeister et al., (2024), referred to as weekly Economic Conditions Indicator (ECI), is suitable for our purposes. This is a composite indicator based on mixed-frequency dynamic factor models covering multiple dimensions of state-level economic activity (mobility, labor market, real activity, expectations measures, financial and households indices). Since the ECI is scaled to match four-quarter growth rates of U.S. real GDP, and we are interested in measuring the economic effects of weather shocks at a monthly frequency, we take the monthly average of the weekly index and label it Monthly Economic Conditions Indicator (MECI) hereafter. Unlike other indicators of economic activity available at the state level on a monthly basis - such as employment or hours worked - the MECI includes a broad set of both standard and non-standard variables (e.g., electricity consumption, oil rig counts, vehicle miles traveled) that allow for a more comprehensive measure, thereby enabling a more thorough analysis of spillover effects. See Bokun et al., (2023) for an overview of state-specific variables for the U.S. [^7] In the online Appendix, we rely on more disaggregated indicators of economic activity (e.g., the number of hours worked in manufacturing) as alternative dependent variables in the GVAR, in place of MECI. The additional results corroborate the baseline ones. However, as expected, using less comprehensive measures of economic activity leads to estimate weaker spillover effects.
@@
Note that we are interested in studying the effects of extreme weather, therefore it is important to assess the presence of seasonality in the MECI. The individual components of the ECI are seasonally adjusted when appropriate (please refer to Baumeister et al., , 2024, for details on seasonal adjustment). Nevertheless, we further check that there are no seasonal effects left in the monthly version of the index by estimating state-specific autoregressive models augmented with month-of-the-year dummy variables. The $F$-tests indicate that for all states, but North Carolina, month-of-the-year dummies are jointly non-significant, and therefore seasonal factors do not affect our results.
@@
Interstate trade flows. In GVAR models, economic interrelationships between geographic entities are most commonly measured using trade flows. To proxy bilateral trade flows between U.S. states, we use data from the latest release of the Commodity Flow Survey (CFS) in 2017. The CFS is jointly elaborated by the Bureau of Transportation Statistics (U.S. Department of Transportation) and the U.S. Census Bureau (U.S. Department of Commerce) every five years as part of the Economic Census. This survey is the primary source of data on domestic freight shipments by establishments in mining, manufacturing, wholesale, and selected retail and services trade industries (namely, electronic shopping and mail-order houses, fuel dealers, publishers) located in the U.S. We use these data to construct weight matrices that summarize network links between states. In particular, to calculate the weight of state $j$ in the network of state $i$, we use the sum of bilateral imports and exports between the two states (i.e., amounts of CFS shipments of goods from $j$ to $i$ and from $i$ to $j$, respectively), as a ratio of total trade of state $i$ (imports into $i$ plus exports from $i$) with all other states in the U.S.
@@
Notes: The left panel shows the classification of U.S. states into nine climate regions, provided by the National Oceanic and Atmospheric Administration (source: https://www.ncei.noaa.gov/access/monitoring/reference-maps/us-climate-regions). Alaska and Hawaii are not considered in this classification but are kept in the analysis. The right panel provides a network visualization of major bilateral trade flows between U.S. states.
When reporting the results, we will aggregate states into climate regions. To this end, we will use the classification adopted by the National Oceanic and Atmospheric Administration (NOAA). Specifically, scientists at the NOAA National Centers for Environmental Information (NCEI) have identified nine climatically consistent regions (Karl and Koss, , 1984). On the left panel of Figure 3 we show these nine climate regions, which are: Northeast (NE), Southeast (SE), South (S), Upper Midwest (UMW), Ohio Valley (OV), Northern Rockies and Plains (NP), Southwest (SW), West (W), and Northwest (NW).
A network representation of the interstate trade flows, instead, is provided in the right panel of Figure 3. Visual clutter is reduced using CFS data to determine, for each state, the three most important trade partners and representing the network as an undirected graph. Nodes are proportional to the size of the local economy, as proxied by state employment in 2017 sourced from the FRED-SD database (Bokun et al., , 2023), and have the same color as the climate region to which they belong, as shown in the left panel of the figure. The network reflects the underlying spatial proximity of states as well as their centrality to the U.S. economy as a whole. For instance, we can observe a cluster of northeastern economies tightly interconnected on the right portion of the graph. Moreover, it is evident that New York, Texas, Florida and California not only represent the largest economies in terms of employment, but are also heavily connected with several other states, thus playing a pivotal role in the entire U.S. economy.
@@
We construct a GVAR model for the U.S. economy by aggregating state-specific autoregressive models (ARX*) that include rest-of-U.S. (RoUS) exogenous variables, along with weather shocks. RoUS aggregate variables are calculated as weighted averages across all other states, using inter-state trade shares as weights. Since RoUS aggregates are treated as weakly exogenous in each ARX* model, estimation can be performed at the state level, thus avoiding overparametrization problems. Let $y_{it}$ denote an indicator of economic activity for state $i$ at time $t$, and $s_{it}$ be our state-level weather-shock proxy. For ease of notation, in the following discussion we consider a model with one lag. The ARX* for state $i$ is then given by:
@@
$\left(\begin{array}[]{c}y_{it}\\ y_{it}^{\ast}\end{array}\right)=\mathbf{W}_{i}\Big{(}y_{1t}\hskip 8.5359pty_{2t}\hskip 8.5359pt\cdots\hskip 8.5359pty_{Nt}\Big{)}^{\prime}=\mathbf{W}_{i}\mathbf{y}_{t}$ (4)
@@
Stacking all ARX* models, we get a country-wide model:
@@
$\hskip-10.0pt\mathbf{G}=\begin{pmatrix}\left(1,-\gamma_{01}\right)\mathbf{W}_{1}\\ \left(1,-\gamma_{02}\right)\mathbf{W}_{2}\\ \vdots\\ \left(1,-\gamma_{0N}\right)\mathbf{W}_{N}\end{pmatrix}\hskip 35.0pt\mathbf{H}=\begin{pmatrix}\left(\beta_{1},\gamma_{11}\right)\mathbf{W}_{1}\\ \left(\beta_{2},\gamma_{12}\right)\mathbf{W}_{2}\\ \vdots\\ \left(\beta_{N},\gamma_{1N}\right)\mathbf{W}_{N}\end{pmatrix}\hskip 35.0pt\boldsymbol{\Theta}=\begin{pmatrix}\theta_{1}&0&\dots&0\\ 0&\theta_{2}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\theta_{N}\end{pmatrix}$ (6)
while $\boldsymbol{\alpha}$, $\mathbf{s}_{t}$ and $\mathbf{u}_{t}$ are vectors of stacked constants, weather shocks and errors, respectively. By inverting the matrix $\mathbf{G}$, we obtain a reduced-form VAR with exogenous variables (VARX) for the entire U.S. economy:
@@ 4 Main results
We estimate our monthly GVAR on data from January 1990 to December 2019. We exclude the subsequent period to prevent our results from being biased by the COVID-19 pandemic. We derive our baseline results using the MECI as a measure of local economic activity. Unit root tests indicate that the level of the MECI is non-stationary or near unit root for a number of states, and stationary for others. Therefore, we transform the MECI indices by taking their first differences and estimate each ARX* model (3) by OLS. [^8] Since not all MECI indices are non-stationary, we cannot represent our U.S. GVAR model as a cointegration model as in Pesaran et al., (2004). Moreover, we have tried estimating the GVAR on the levels of the MECI, treating them as stationary, but the model turns out to have explosive roots in this case. Reliance on first differences of MECI implies that our model does not have an error correction form. For the same reason, we cannot rely on the usual likelihood ratio (LR) test for weak exogeneity of $y_{it}^{*}$ used in the cointegrated GVAR literature, which tests the significance of loading coefficients in error-correction country-specific models (see, e.g., Dées et al., 2007a ). However, we perform Granger causality tests on $y_{it}$ and $y_{it}^{*}$ in each state, separately. The tests indicate that, for the vast majority of states, including the ones with the largest economies, the domestic variable $y_{it}$ does not Granger cause the rest-of-U.S. variable $y_{it}^{*}$, while $y_{it}^{*}$ Granger causes $y_{it}$, as expected. Details in the online Appendix. We rely on a GVAR model with 2 lags of first-differenced MECI, as suggested by minimization of the Bayesian information criterion, which selects at most 2 lags for almost all states. [^9] For robustness, we have also estimated the model using, alternatively, lag orders of 1 and 3. The results remain very similar in both cases. Furthermore, in consideration of the monthly frequency of the data, we have also estimated the model with 12 lags. The results are qualitatively similar also in this case.
@@
We assess the direct and indirect macroeconomic effects of weather shocks by means of impulse response functions (IRFs) estimated on a sample that runs from April 1990 through December 2019, for a total of 357 observations (i.e., after taking first differences of MECI and considering two lags). While we carry out the impulse-response analysis at the state level, in what follows, to summarize results, we aggregate the estimated individual IRFs into those for nine U.S. climate regions, as previously shown in Figure 3. Note that while Alaska and Hawaii are not considered in this classification, we do retain them in the GVAR model.
@@
The impulse response functions are obtained by simulation, i.e., by projecting h steps ahead the fitted indicator of economic activity, with and without the impulse of the weather shock, and taking the difference between the two projected paths. When simulating a weather shock, we assume that all states in a region are contemporaneously hit, and all counties within each state. As a matter of fact, Figure 2 highlights that major weather events such as the Storm of the Century in 1993, the Blizzard of 1996, Hurricane Katrina in 2005 and the Super Outbreak in 2011 have triggered emergency declarations in a very large portion of the U.S. This is thus an extremely adverse but realistic scenario, that allows us to interpret the results as “upper bounds” (in absolute value) of cross-regional spillover effects.
4.1 The impacts of weather shocks on economic activity
@@
We assess the responses of MECI to a region-wide weather shock in Figures 4-6. While the GVAR model is estimated using the first differences of MECI, we cumulate the IRFs to obtain the responses in terms of level of the index, which is expressed as a year-on-year percent growth rate (see section 2.2). Thus, an IRF value of -1 means that a shock reduces economic activity by about 1% compared to one year before. As discussed in the Appendix, where we also report the IRFs for the first differences of the index ($\Delta\text{MECI}$), the impact on month-on-month growth rates generally vanishes within a year after a shock. Figure 4 shows the IRFs for a subset of regions (namely, NE, SE, S, and W), while Figure 5 reports the IRFs for all regions in the month when the shock occurs, and Figure 6 reports the IRFs for all regions one year after the shock, which seems a reasonable time frame to assess the magnitude of macroeconomic effects induced by weather disasters. We compute mean IRFs and include both the 80% (i.e., based on the 10th and 90th percentiles) and 95% (i.e., based on the 2.5th and 97.5th percentiles) confidence intervals using a bootstrap procedure for GVAR models (see Dées et al., 2007a, ; Dées et al., 2007b, , for details). [^10] The bootstrap procedure is non-parametric, so we do not need to make any assumption on the distribution of errors $u_{it}$ in equation (3). In the Appendix, we also report results for the 5th and the 95th percentiles (i.e., the 90% confidence interval). Since the results are qualitatively comparable, in what follows we comment on the results focusing on the 80% confidence interval.
Remarkably, the responses of economic activity are generally negative, both within the shocked region and in regions not directly affected by the weather shock. Except for Western regions (W, NW, and SW), all regions generate significant spillovers. The cumulative IRFs converge within approximately one year after the shock, which is consistent with monthly economic activity being affected for periods lasting up to one year.
@@
Notes: Responses of the Monthly Economic Conditions Indicator (MECI) to a weather shock hitting 100% of counties in the shocked region. The horizontal axis measures months, the vertical axis the responses in percentage points. Blue solid lines indicate bootstrap means and light-red (dark-red) areas the corresponding 2.5th and 97.5th (10th and 90th) percentiles. Four climate regions are considered: Northeast (NE), Southeast (SE), South (S), and West (W). In the titles, the first region indicates where the shock originates, while the second one is where it propagates.
As already said, Figure 5 and Figure 6 report the mean bootstrap responses along with confidence intervals on the month of the shock and one year after the shock, respectively. The largest spillovers (in absolute values) are generated by the South (S), Southeast (SE) and Ohio Valley (OV) regions. Concerning the magnitude of the effects, the mean responses after one year are generally in the range from -1.8 to 0. This indicates, approximately, a drop of up to 1.8% in year-on-year growth rates of economic activity, in response to the region-wide weather shock. In interpreting the results, it should be kept in mind that we are reporting responses to an upper-bound shock hitting 100% of counties in all states within a region. Of course, extreme events typically hit only a fraction of counties. On average, in months when weather disasters strike, about 10% of counties in a region are affected, based on FEMA emergency declarations. To assess the responses to a shock hitting 10% of counties, as an example, the reported IRFs should be divided by 10. Thus, for instance, a shock of this size occurring in the South can reduce growth by 0.09%-0.18% in other regions, approximately.
Further results and robustness checks are presented in Section 5. Moreover, in the Appendix we also replicate the analysis for different categories of weather-related disasters.
Figure 5: IRFs of Monthly Economic Conditions Indicators, impact on month of the shock
Notes: The figure displays IRF bootstrap means, along with the corresponding 2.5th and 97.5th (10th and 90th) percentiles, represented as points (bars). The horizontal axis specifies the regions responding to shocks, while the title identifies the origin of the shocks.
@@
Notes: See notes to Figure 5.
4.2 The role of economic linkages and spatial proximity
The GVAR model allows us to estimate spillover effects taking into account the economic interdependence between states. In this section, we assess the importance of explicitly considering economic linkages through interstate trade flows. To this aim, we re-estimate the GVAR model using alternative weight matrices $\mathbf{W}_{i}$ based on spatial adjacency.